# An Exposition of Pathfinding Strategies Within Lightning Network Clients

### v1

@@ VI-B CLN
for the channel $e$. Similar to its application in LND as observed in (9) with $\operatorname{\Delta^{\text{CLTV}}_{e}}$, the $\mathrm{riskfactor}$ is used here in conjunction with $\operatorname{\Delta^{\text{CLTV}}_{e}}$ to regulate its impact. The default value for the $\mathrm{riskfactor}$ is 10. Essentially, in (17), the percentage of yearly expense incurred by funds trapped within the channel is considered. With 52,596 blocks per year, the yearly cost per block equates to the $\frac{1}{5,259,600}$.
Additionally, CLN introduces $\mathrm{capacityBias}_{e}$ to penalize channels where a substantial portion of the channel’s total capacity is being utilised. The formulation of $\mathrm{capacityBias}_{e}$ is based on modeling the channel-wise success probability under the assumption of uniform liquidity distribution mentioned in [14]. It is given by,
$\mathrm{capacityBias}_{e}=-\log\left(\frac{\mathrm{cap}_{e}+1-\mathrm{amt}_{e}}{\mathrm{cap}_{e}+1}\right).$ (18)
(18) is similar to (6) with $P_{e}=\frac{\mathrm{cap}_{e}+1-\mathrm{amt}_{e}}{\mathrm{cap}_{e}+1}$.
Analysis of Equation 17 illustrates a notable emphasis on $\operatorname{\Delta^{\text{CLTV}}_{e}}$through the selection of higher value for the $\mathrm{riskfactor}$. Consequently, CLN is inclined to favor paths characterized by lower timelock.
VI-C LDK
LDK incorporates design elements such as fees and HTLC minimum into its weight function, along with penalty. The weight function for using the channel $e$ in LDK is as given below,
$\mathrm{weight(e)}=\max(\mathrm{fee}_{e},\ \mathrm{pathHtlcMin}_{e})+\mathrm{penalty}_{e}.$ (19)
In routing, payments must meet the minimum HTLC limit set by each channel that is used as described in Section IV. Therefore, to utilize the channel by meeting the HTLC minimum, the payment source might have to pay an additional amount along with the actual payment amount. This additional amount is denoted as $\mathrm{pathHtlcMin}_{e}$. $\mathrm{HtlcMin}_{e}$ in the equation below is obtained from the channel update message [48].
$\mathrm{pathHtlcMin}_{e}$ $=\mathrm{HtlcMin}_{e}\cdot(1+\mathrm{feeRate}_{e})$ (20)
$\quad+\mathrm{BaseFee}_{e}.$
$\mathrm{penalty}{{}_{e}}$ in the channel weight is the sum of base penalty, anti probing penalty and total liquidity penalty, expressed as:
@@ VI-C LDK
$\quad+\mathrm{liquidityPenalty}_{e}+\mathrm{historicPenalty}_{e}.$
Base penalty comprises of a fixed penalty $\mathrm{penaltyBase}$, which is applied to each channel has default value 500 msat, and $\mathrm{baseMultiplier}$, which is used with amount in the channel, with default value 8,192 msat. The purpose of using $\mathrm{baseMultiplier}$ with amount is to avoid having fees dominate the channel weight for large amount.
$\mathrm{basePenalty}_{e}=\mathrm{penaltyBase}+\frac{\mathrm{baseMultiplier}\cdot\mathrm{amt}_{e}}{2^{30}}.$ (22)
@@
In $\mathrm{liquidityPenalty}{{}_{e}}$ calculation, a multiplier $\mathrm{LM}$ is used, with the negative logarithm of the channel-wise success probability $P_{e}$ for a payment. The preset value of this multiplier is 30,000 msat. Additionally, an $\mathrm{amtMultiplier}_{e}$ is used in $\mathrm{liquidityPenalty}{{}_{e}}$ calculation. The penalty is influenced by the latest estimate of the liquidity in the channel.
@@
$\quad\cdot(\mathrm{LM}+\mathrm{amtMultiplier}_{e}).$
$\mathrm{historicPenalty}{{}_{e}}$ is similar to $\mathrm{liquidityPenalty}{{}_{e}}$. However, in this case, instead of solely relying on the latest estimate of the liquidity available in the channel, the calculation of the success probability is based on the historical estimated liquidity available in the channel. A multiplier $\mathrm{HM}$ is also utilized alongside the negative logarithm of the channel’s success probability $P_{e}$ for the payment. This calculation is based on the historical estimates of the channel’s available liquidity. It takes the default value 10,000 msat.
$\mathrm{historicPenalty}_{e}$ $=-\log_{10}(P_{e})$ (25)
@@
$\mathrm{amtMultiplier}_{e}$ in Equation 24 and Equation 25 is given by,
$\mathrm{amtMultiplier}_{e}=\frac{\mathrm{Multiplier}\cdot\mathrm{amt}_{e}}{2^{20}}.$ (26)
[…]
A channel’s success probability $P_{e}$ is calculated using upper and lower liquidity bounds $\mathrm{UB}_{e}$ and $\mathrm{LB}_{e}$ that are determined using the past unsuccessful and successful payments based on the formula
$P_{e}=\frac{\mathrm{UB}_{e}-\mathrm{amt}_{e}}{\mathrm{UB}_{e}-\mathrm{LB}_{e}}.$ (27)
This calculation, similar to (18), also based on modeling the success probability under uniform liquidity distribution assumption as described in [14].
@@ VI-D1 Case 1: Heuristics ratios
$\mathrm{weight(e)}=(\mathrm{fee}_{e}+\mathrm{HopCost}_{e})\cdot\mathrm{factor}_{e}$ (28)
where $\mathrm{factor}_{e}$ is given by,
@@
$\quad+0.99998\cdot\frac{\min(\max(\min D,v),\max D)-\min D}{\max D-\min D}$
@@ VI-D3 Case 3: Heuristics constants (with logarithm)
In routing, as mentioned in Section IV, there is a constant risk of funds being locked in HTLC. This risk is represented by the heuristic constant in Eclair, denoted as $\mathrm{lockedFundsRisk}$, with a preset value of $10^{-8}$. Therefore, $\mathrm{RiskCost}_{e}$ is calculated to manage the influence of timelock constraints.
@@
Failure cost represents a virtual cost associated with failed payment attempt, similar to the use of attempt cost in LND (11).
@@
$\quad+\mathrm{amt}_{e}\cdot\mathrm{FailureCostRate}$
@@
The default values of $\mathrm{BaseHopCost}$ and $\mathrm{HopCostRate}$ are both zero. Fine-tuning $\mathrm{HopCost}_{e}$ can be considered to optimize for different performance metrics.
The fees associated with per-hop charges and failed attempts are not actually spent. They function as incentives to favor shorter paths or routes with greater probability of success.
@@ VII Evaluation
Our simulation model is developed using Python 3.11 for all experiments. The source code is available on GitHub at: https://github.com/sindurasaraswathi/LN-PathFinding.
The Lightning Network graph is constructed from the dataset [40], and all graph-related operations are performed using the Networkx library [49]. The utilized Lightning Network snapshot [40] comprises 13,129 nodes and 57,773 channels, providing channel-specific information such as short channel ID, channel ID, capacity, base fee, proportional fee, minimum and maximum HTLC, and CLTV delta.
All relevant information required for pathfinding and used in weight functions as design elements is included as attributes of the channel in the generated graph. Channel age is extracted using the short channel ID. We constructed two LN graphs. The first graph is built by uniformly choosing channel balances between zero and the channel capacity. The second graph is constructed by sampling channel balances from a bimodal distribution $\mathrm{P}(x)\sim e^{\frac{\mathrm{-x}}{\mathrm{s}}}+e^{\frac{\mathrm{x-cap}_{e}}{\mathrm{s}}}$, where $\mathrm{s}$ is selected as a fixed fraction of the channel capacity, specifically $\mathrm{s}=\frac{\mathrm{cap}_{e}}{10}$. We refer to the first and second graphs as the uniform distribution graph and the bimodal distribution graph, respectively, throughout this section.
@@
Understanding the routing behavior of nodes across these different connectivity levels is also crucial. Hence, we classify the nodes into well-connected, fairly connected, and poorly connected categories based on the total capacity across all channels originating from a node and the number of channels. Specifically, nodes with a total capacity greater than or equal to $10^{6}$ sats and more than 5 channels are categorized as well-connected, those with a total capacity falling between $10^{4}$ and $10^{6}$ sats and more than 5 channels are classified as fairly connected, and nodes with 5 or fewer channels are designated as poorly connected, regardless of their total channel capacity. The statistics on node connectivity, including the number of nodes and channels, as well as the channel capacity for each node category, are presented in Table II.
@@
We conducted several experiments to compare the empirical performance of the four Lightning clients and their variations. Initially, using the uniform distribution graph, simulations were performed for 10,000 transactions, with source and receiver nodes randomly selected from all nodes in the graph. Transaction amounts were uniformly chosen between 1 and the minimum of the maximum values of all outgoing channel balances of the source node and all local incoming balances of the receiver node. By using this upper bound, we were able to select payment amounts that are practically feasible for the source node to send and for the receiver node to receive.
@@
We used the _$\_dijkstra$ method from the Networkx library to execute the Dijkstra’s algorithm for CLN and LDK, and adapted _$\_dijkstra$ for LND as discussed in Section V, supplying the method with an appropriate weight function and side constraints. For finding Yen’s K shortest paths in the case of Eclair, we adapted the $shortest\_simple\_paths$ method from the Networkx library. Upon determining a path for a given lightning variant and transaction, our model simulated the routing of the payment amount through the determined path. The routing was regarded a failure if no acceptable path could be found or if any of the intermediaries nodes had insufficient balances.
@@
In our experiments, we studied and compared a total of seven variants of LN clients: LND1, LND2, CLN, LDK, Eclair__\__case1, Eclair__\__case2, and Eclair__\__case3. LND1 and LND2 represent the LND implementation with Apriori and Bimodal estimators, respectively, for channel-wise success probability calculation. Similarly, Eclair_case1, Eclair_case2, and Eclair__\__case3 denote three distinct cases of Eclair: heuritics ratios, heuristics constants (without logarithm) and heuristics constants (with logarithm), respectively, as discussed in Section VI.
Figure 2: Rate of success of different lightning clients across transaction amount bins
(a) LND1
(b) LND2
(c) CLN
(d) LDK
(e) Eclair__\__case1
(f) Eclair__\__case2
(g) Eclair__\__case3
Figure 3: Fee distribution across transaction amount bins
Figure 4: Distribution of path lengths across Lightning Network clients
We investigated the behavior of each variant of the LN clients across various transaction amount bins, spanning the range from 1 to $10^{8}$. From Figure 2, it is evident that LND2 exhibits the highest success rate and excels particularly with larger payment amounts. Conversely, LND1 demonstrates the lowest success rates in the mid-range amount bins. CLN’s success rate remains competitive for larger amounts, while LDK performs well by delivering the highest success rates for small and lower mid-range amounts; however, its success rate experiences a significant decline for larger amounts. Eclair_case1 and Eclair_case2 demonstrate similar performance levels, with higher success rates observed for larger payment amounts compared to LDK and CLN. Meanwhile Eclair_case3 exhibits higher success rates than Eclair_case1 and Eclair_case2 and its performance is closer to LND2, making it a competitive performer in terms of success rate.

### v2

@@ VI-B Core Lightning (CLN)
for the channel $e$. Similar to its application in LND as observed in (13) with $\operatorname{\Delta^{\text{CLTV}}_{e}}$, a constant $\mathrm{riskfactor}$ is used here in conjunction with $\operatorname{\Delta^{\text{CLTV}}_{e}}$ to regulate its impact. The default value for $\mathrm{riskfactor}$ is $10$. The denominator in (20) is motivated from a yearly expense incurred by funds locked within a channel, and uses the constant $\mathrm{BlockPerYear}=52,596$ due to the expected number of blocks per year in the Bitcoin network.
Additionally, CLN introduces $\mathrm{capacityBias}_{e}$ to penalize channels where a substantial portion of the channel’s total capacity is being utilized. The formulation of $\mathrm{capacityBias}_{e}$ can be derived from a modeling of the path success probability $P_{e}$ with a negative logarithm penalization (cf. (7)) and the assumption of a uniform liquidity distribution similar to (8), and is given by
$\mathrm{capacityBias}_{e}=-\log\left(\frac{\mathrm{cap}_{e}+1-\mathrm{amt}_{e}}{\mathrm{cap}_{e}+1}\right),$ (21)
where $\mathrm{amt}_{e}$ is the channel payment amount and $\mathrm{cap}_{e}$ is the channel capacity.
Overall, the breakdown of (20) illustrates a notable emphasis of CLN’s pathfinding on $\operatorname{\Delta^{\text{CLTV}}_{e}}$ through the selection of higher value for the $\mathrm{riskfactor}$. Consequently, CLN is inclined to favor paths characterized by lower timelock. CLN uses this weight function (20) in Algorithm 1 and constraint $\sum_{e\in p}1\leq 10$ is validated after pathfinding.
VI-C Lightning Development Kit (LDK)
LDK’s pathfinding incorporates the design elements fees, success probability, HTLC minimum, HTLC maximum into its weight function. The resulting channel weight function of LDK can be defined via
$\mathrm{weight(e)}=\max(\mathrm{fee}_{e},\ \mathrm{pathHtlcMin}_{e})+\mathrm{penalty}_{e},$ (22)
where $\mathrm{fee}_{e}$ is the total fee (3), $\mathrm{penalty}_{e}$ a multifaceted term discussed below, and $\mathrm{pathHtlcMin}_{e}$ a prospective fee cost of related to a payment of size of the minimum HTLC. In particular, payments must meet the minimum HTLC limit set by each channel that is used as described in Section IV. Therefore, to utilize the channel by meeting the HTLC minimum, the payment source might have to pay an additional amount along with the actual payment amount, leading to the term
$\mathrm{pathHtlcMin}_{e}$ $=\mathrm{HtlcMin}_{e}\cdot(1+\mathrm{f}_{e})+\mathrm{b}_{e},$ (23)
where $\mathrm{HtlcMin}_{e}$ is the channel-specific HTLC minimum parameter, $\mathrm{b}_{e}$ the base fee and $\mathrm{f}_{e}$ the proportional fee charged by the channel, which are all obtained from the channel update message [8].
The term $\mathrm{penalty}{{}_{e}}$ in the channel weight is the sum of four terms related to a base penalty, an anti-probing penalty, a liquidity penalty and a penalty based on a historical estimate, such that
@@ VI-C Lightning Development Kit (LDK)
$\quad+\mathrm{liquidityPenalty}_{e}+\mathrm{historicPenalty}_{e},$
where the base penalty
$\mathrm{basePenalty}_{e}=\mathrm{penaltyBase}+\frac{\mathrm{baseMultiplier}\cdot\mathrm{amt}_{e}}{2^{30}}$ (25)
consists of a fixed penalty $\mathrm{penaltyBase}$ with a default value of $500$ msat and the constant $\mathrm{baseMultiplier}$, which is multiplied to the payment amount $\mathrm{amt}_{e}$ of the channel and whose default value is $\mathrm{baseMultiplier}=8,192\,\mathrm{msat}$.
@@
The liquidity penalty, which incorporates success probability modeling, is defined as
@@
$\cdot(\mathrm{LM}+\mathrm{amtMultiplier}_{e}),$
where $\mathrm{LM}=30,000\,\mathrm{msat}$ is a constant multiplier, $\mathrm{amtMultiplier}_{e}$ is defined below in (29), and $P_{e}$ is a channel success probability estimate (see discussion of $P_{e}$ at the end of this subsection), based on the latest estimate of channel liquidity.
$\mathrm{historicPenalty}{{}_{e}}$ is similar to $\mathrm{liquidityPenalty}{{}_{e}}$. However, in this case, instead of solely relying on the latest estimate of the liquidity available in the channel, the calculation of the success probability is based on the historical estimated liquidity available in the channel. A constant multiplier $\mathrm{HM}$ is also utilized alongside the negative logarithm of the channel’s success probability $\mathrm{hP_{e}}$ for the payment. The calculation of $\mathrm{hP_{e}}$ is also performed using $P_{e}$, as discussed at the end of this section, based on historical estimates of the channel’s available liquidity, rather than the most recent liquidity estimates, as in the case of $P_{e}$ in $\mathrm{liquidityPenalty}_{e}$. $\mathrm{historicPenalty}_{e}$, with the constant multiplier $\mathrm{HM}=10,000\,\mathrm{msat}$, is defined as
$\mathrm{historicPenalty}_{e}$ $=-\log_{10}\mathrm{(hP_{e})}$ (28)
@@
$\mathrm{amtMultiplier}_{e}$ in (27) and (28) is given by,
$\mathrm{amtMultiplier}_{e}=\frac{\mathrm{Multiplier}\cdot\mathrm{amt}_{e}}{2^{20}},$ (29)
where $\mathrm{Multiplier}=\mathrm{liquidityPenaltyMultiplier}=192\,\mathrm{msat}$ for $\mathrm{liquidityPenalty}{{}_{e}}$ and $\mathrm{Multiplier}=\mathrm{historicPenaltyMultiplier}=64\,\mathrm{msat}$ for $\mathrm{historicPenalty}{{}_{e}}$, respectively.
Within LDK, there are two different options for estimating the channel success probability $P_{e}$: the first assumes a uniform channel liquidity distribution, and the second assumes a bimodal liquidity distribution. Under the uniform liquidity assumption, $P_{e}$ is calculated using upper and lower liquidity bounds $\mathrm{UB}_{e}$ and $\mathrm{LB}_{e}$ that are determined using the past unsuccessful and successful payments based on the formula
$P_{e}=\frac{\mathrm{UB}_{e}-\mathrm{amt}_{e}}{\mathrm{UB}_{e}-\mathrm{LB}_{e}},$ (30)
see also (9).
As a second option, LDK can assume a bimodal liquidity distribution given by $P(x)\sim\Big{(}\mathrm{x-\frac{cap_{e}}{2}}\Big{)}^{2}$, where channel liquidities are concentrated at the edges. Under this assumption, $P_{e}$ is computed as the integral under the curve from the payment amount in the channel to the channel’s upper bound, divided by the integral under the curve from the lower bound to the upper bound of the channel liquidity. This is expressed as:
$P_{e}$ $=\frac{\int_{\mathrm{amt}_{e}}^{\mathrm{UB}_{e}}\mathrm{P(x)}\,\mathrm{dx}}{\int_{\mathrm{LB}_{e}}^{\mathrm{UB}_{e}}\mathrm{P(x)\,dx}}$ (31)
Therefore, we obtain
$P_{e}$ $=\frac{\left((\mathrm{UB}_{e}-\frac{\mathrm{cap}_{e}}{2})^{3}-(\mathrm{amt}_{e}-\frac{\mathrm{cap}_{e}}{2})^{3}\right)\cdot 64\cdot 1024^{3}+1}{\left((\mathrm{UB}_{e}-\frac{\mathrm{cap}_{e}}{2})^{3}-(\mathrm{LB}_{e}-\frac{\mathrm{cap}_{e}}{2})^{3}\right)\cdot 64\cdot 1024^{3}+1}.$ (32)
@@
LDK uses the weight function (22) in Algorithm 1. Similar to LND, LDK also imposes side constraints during pathfinding, as outlined in Table I. These include: the timelock constraint $\sum_{e\in p}\operatorname{\Delta^{\text{CLTV}}_{e}}\leq 1008$, the probability constraint $-\sum_{e\in p}\log(P_{e})\leq-\log(0.01)$, the fee constraint $\sum_{e\in p}\mathrm{fee}_{e}\leq\mathrm{(1\%\ of\ payment\ amount+50\ sats)}$, and the path length constraint $\sum_{e\in p}1\leq\mathrm{19}$.
@@ VI-D1 Case 1: Heuristics ratios
$\mathrm{weight(e)}=(\mathrm{fee}_{e}+\mathrm{HopCost}_{e})\cdot\mathrm{factor}_{e},$ (33)
where $\mathrm{factor}_{e}$ is given by
@@
$\quad+0.99998\cdot\frac{\min(\max(\min D,v),\max D)-\min D}{\max D-\min D},$
@@ VI-D3 Case 3: Heuristics constants (with logarithm)
In routing, as mentioned in Section IV, there is a constant risk of funds being locked in HTLC. This risk is represented by the heuristic constant in Eclair, denoted as $\mathrm{lockedFundsRisk}$, with a preset value of $10^{-8}$. Therefore, $\mathrm{RiskCost}_{e}$ is calculated to manage the influence of timelock constraints and is given as:
@@
Failure cost represents a virtual cost associated with failed payment attempt, similar to the use of attempt cost in LND (15), defined as:
@@
$\quad+\mathrm{amt}_{e}\cdot\mathrm{FailureCostRate},$
@@
Similarly, $\mathrm{HopCost}_{e}$, a virtual cost to optimize the path length, is defined as:
@@
The default values of $\mathrm{BaseHopCost}$ and $\mathrm{HopCostRate}$ are both zero. A higher $\mathrm{HopCost}_{e}$ strongly penalizes longer paths. Fine-tuning $\mathrm{HopCost}_{e}$ can be considered to optimize for path length. The fees associated with per-hop charges and failed attempts are not actually spent. They function as incentives to favor shorter paths with greater probability of success.
Eclair integrates its weight function into Yen’s $K$-shortest path algorithm for pathfinding, with the default number of paths set to $K=3$. After finding paths, it validates constraints such as the timelock limit $\sum_{e\in p}\operatorname{\Delta^{\text{CLTV}}_{e}}\leq\mathrm{2016}$, the fee limit $\sum_{e\in p}\mathrm{fee}_{e}\leq\mathrm{User\ Defined\ Max\_Fee\_Limit}$, and the path length limit $\sum_{e\in p}1\leq\mathrm{20}$.
@@ VII Evaluation
VII-A Simulation Model
Our simulation model is developed using Python 3.11 for all experiments. The source code is available on GitHub at: https://github.com/sindurasaraswathi/LN-PathFinding. The Lightning Network graph is constructed from the dataset [44], and all graph-related operations are performed using the Networkx library [52]. The utilized Lightning Network snapshot [44] comprises 13,129 nodes and 57,773 bi-directional channels, providing channel-specific information such as short channel ID, channel ID, capacity, base fee, proportional fee, minimum and maximum HTLC, and CLTV delta.
All relevant information required for pathfinding and used in weight functions is included as attributes of the channel in the generated graph. Channel age is extracted using the short channel ID. Consistent with the available information of an arbitrary LN node operator who is not engaged in balance probing, the dataset [44] does not contain information about channel balances, which is only known to the two participating nodes of channel. To simulate the channel balances, we use two random models for the balances. The first model is built by uniformly choosing channel balances between zero and the channel capacity. The second graph model is constructed by sampling channel balances from a bimodal distribution with density $\mathrm{P}(x)$ proportional to
$\mathrm{P}(x)\sim e^{\frac{\mathrm{-x}}{\mathrm{s}}}+e^{\frac{\mathrm{x-cap}_{e}}{\mathrm{s}}}$ (42)
on $[0,\mathrm{cap}_{e}]$, where $\mathrm{s}$ is selected as a fixed fraction of the channel capacity, specifically $\mathrm{s}=\frac{\mathrm{cap}_{e}}{10}$. We refer to the first and second graphs as the uniform distribution graph and the bimodal distribution graph, respectively, throughout this section.
@@ VII-A Simulation Model
Understanding the routing behavior of nodes across these different connectivity levels is also crucial. Hence, we distinguish well-connected, fairly connected, and poorly connected nodes based on the total capacity across all channels originating from a node and the number of channels a node is participating in. Specifically, nodes with a total capacity greater than or equal to $10^{6}$ sats and more than 5 channels are called well-connected, those with a total capacity falling between $10^{4}$ and $10^{6}$ sats and more than 5 channels are called fairly connected, and nodes with $5$ or fewer channels are designated as poorly connected, regardless of their total channel capacity. The statistics on node connectivity, including the number of nodes and channels, as well as the channel capacity for each node category, are presented in Table II.
@@
We conduct several experiments to compare the empirical performance of the four Lightning clients and their variations. Initially, using the uniform distribution graph, simulations are performed for 10,000 transactions, with source and receiver nodes randomly selected from all nodes in the graph. Transaction amounts are uniformly chosen between $1$ and the minimum of the maximum values of all outgoing channel balances of the source node and all local incoming balances of the receiver node. By using this upper bound, we can select payment amounts that are practically feasible for the source node to send and for the receiver node to receive.
@@
We use the _$\_dijkstra$ method from the NetworkX library to execute the Dijkstra’s algorithm for CLN and LDK, and adapted _$\_dijkstra$ method for LND as discussed in Section V, supplying the method with an appropriate weight function. For finding Yen’s K shortest paths in the case of Eclair, we adapt the $shortest\_simple\_paths$ method from the NetworkX library. Upon determining a path for a given Lightning client and transaction, our model simulates the routing of the payment amount through the determined path. The routing is regarded a failure if no acceptable path could be found or if any of the intermediary nodes has insufficient balances.
@@
In our experiments, we study and compare a total of eight variants of LN clients: LND-ap, LND-bm, CLN, LDK-un, LDK-bm, Eclair1, Eclair2, and Eclair3. LND-ap and LND-bm represent the LND implementation with $Apriori$ (16) and $Bimodal$ (19) estimators, respectively, for channel-wise success probability calculation. The LDK implementation with $Uniform$ (30) and $Bimodal$ (32) estimators for channel-wise success probability is represented as LDK-un and LDK-bm, respectively. Similarly, Eclair1, Eclair2, and Eclair3 denote three distinct cases of Eclair: heuristics ratios, heuristics constants (without logarithm) and heuristics constants (with logarithm), respectively, as discussed in Section VI. Building on the existing LND implementation, we propose and study a new weight function by assuming a uniform liquidity distribution for calculating channel-wise success probability, defined as
$P_{e}=\frac{\mathrm{cap_{e}-amt_{e}}}{\mathrm{cap_{e}}}.$ (43)
The expression (43) is used to compute the path success probability, which is subsequently incorporated into the LND weight function (13). We refer to this implementation as LND-un. In the experiments, we use the default values for all constants in the weight function of each LN client variant, as mentioned in Section VI.
VII-B Simulation Results
VII-B1 LN Graph with Uniform Balance Distribution
We investigate the behavior of each variant of the LN clients across various transaction amount bins, spanning the range from 1 to $10^{8}$. As a first experiment, whose results are reported in Tables III, IV and V, we sample channel balances from the uniform distribution.
We observe in Table III that for the uniform distribution graph, LND-un exhibits the highest success rate across most of the amount bins compared to all other existing LN client variants. Among the eight existing LN client variants, Eclair3 demonstrates high success rates. LND-ap and LND-bm show similar performance levels, with LND-bm slightly outperforming LND-ap in success rates. CLN’s success rate remains competitive for larger amounts and outperforms LND-ap, while LDK-un and LDK-bm perform moderately for small and lower mid-range amounts; however, their success rates experience a significant decline for larger amounts. Eclair1 and Eclair2 demonstrate similar performance levels, with higher success rates observed for larger payment amounts compared to CLN, LDK variants, and LND variants.